OpenStudy (jazzyluvsu123):

1 and 3/10 as percent is 130% right @jdoe0001

OpenStudy (jdoe0001):

\(\bf 1\frac{3}{10}\cdot 100\implies \cfrac{1\cdot 10+3}{\cancel{10}}\cdot \cancel{100}\implies 13\cdot 10\)

OpenStudy (jazzyluvsu123):

so answer would be 131.0% ? @jdoe0001

OpenStudy (jdoe0001):

yes... but that's how you'd get the %, you'd multiply by 100 and in the case of a fraction, you'd make it an non-mixed fraction first, then multiply by 100

OpenStudy (jazzyluvsu123):

ok thanks @jdoe0001